# Towards Motion Compensation in Autonomous Robotic Subretinal Injections

### v1

@@ V DISCUSSION AND FUTURE WORK
Our results demonstrate that our robotic system is capable of matching the sinusoidal motion of the eye in the Z-axis. For the 100 and 50 $\mu m$ amplitude motions, our robot is able to match their movement. But for the 25 $\mu m$ amplitude motion, we see that the 0.2 mm/s end-effector velocity does not match the motion of the linear stage. This could be due to limitations of our robots ability to move at low velocities. This problem is solved by increasing the applied velocity to 0.3 mm/s.
One of the main limitations observed is a phase shift between the needle’s motion and the retinal tissue. This latency arises from our position correction method, which relies on comparing each incoming B5-scan with the previously processed scan. Consequently, the system continously adjusts the robot’s position based on the prior state of the retina rather than its real-time position. As a result, each corrective movement is slightly delayed, leading to the phase lag seen in the graphs in Fig. 5 and in the OCT images in Fig. 1. This lag affects the precision of needle tracking. It suggests that our system may struggle to maintain synchronization with rapidly moving retinal tissue.
A secondary issue noted in the results is the gradual vertical drift in the needle’s position relative to the retina. Over time, the needle tip deviates from its starting position, shifting either upwards or downwards. This phenomenon is especially prevalent in the 25 $\mu m$ amplitude retina motion examples (Fig. 5 (c) and (d)). During our secondary experiments with subretinal injections the vertical shift also caused the needle to move away from the target depth, leading to water being injected into the vitreous.
@@
Figure 8: OCT B-scan image captured at the 20-second mark of the insertion shown in Figure 6. Yellow circle marks the needle tip, the needle is highlighted in red, and the ILM layer of the retina is highlighted in green.
Another observation is the altered retina motion associated with the needle insertion. Looking at Fig. 6, leaves the tissue by around 100 $\mu m$ when the needle and linear stage motions are out of sync. However, when looking at the OCT scans, we realize that the needle tip remains in the retina and the horizontal shift is much less significant. When the needle penetrates the retina, it likely induces a surface adhesion and resistance within the tissue, restricting free movement. This effect is particularly evident during downward movements, where the needle’s presence within the tissue exerts a tethering force that impedes motion. In Fig. 8, we show an OCT B-scan at the 20-second mark where the needle tip is still in contact with the retina. According to the graph, the needle should be outside of the retinal tissue at this point. Therefore, when tool-tissue interactions are present, relying solely on the linear stage controller’s position, which does not fully represent the tool-retina positional relationship.
@@ VI CONCLUSION
In this work, we presented our initial method and experiments for extending our previously developed autonomous robotic subretinal injection pipeline with motion compensation capabilities along the $Z$-axis. Based on real-time acquisition and segmentation of small 3D volumes, called B5-scans, we calculate the depth of the ILM layer and compare it with the previously computed values. We then move the robot end-effector to keep the needle tip in the same position relative to the retina.
Although our method shows the potential of our robotic system to compensate for simulated motion similar to motion caused by respiration or circulation,some aspects should be further examined, such as interactions of moving tissue and surgical tools and the implementation of predictive methods for estimating retinal motion for more accurate compensation. We plan on working on this topic in the future.

### v2

@@ V DISCUSSION AND FUTURE WORK
Our results demonstrate that our robotic system can match the sinusoidal motion of the eye on the Z-axis. For the 100 and 50 µm amplitude motions, our robot can match their movement. But for the 25 µm amplitude motion, we see that the 0.2 mm/s end-effector velocity does not match the motion of the linear stage. This could be due to limitations of our robot’s ability to move at low velocities. This problem is solved by increasing the applied velocity to 0.3 mm/s.
One of the main limitations observed is a phase shift between the needle’s motion and the retinal tissue. This latency arises from our position correction method, which compares each incoming B5-scan with the previously processed scan. Consequently, the system continuously adjusts the robot’s position based on the prior state of the retina rather than its real time position. As a result, each corrective movement is slightly delayed, leading to the phase lag seen in the graphs in Fig. 5 and the OCT images in Fig. 1. This lag affects the precision of needle tracking. It suggests that our system may struggle to maintain synchronization with rapidly moving retinal tissue.
A secondary issue noted in the results is the gradual vertical drift in the needle’s position relative to the retina. Over time, the needle tip deviates from its starting position, shifting either upwards or downwards. This phenomenon is especially prevalent in the 25 µm amplitude retina motion examples (Fig. 5 (c) and (d)). Additionally, there are examples in our results where the needle velocity is zero (position line is horizontal) or changes direction multiple times in a short time. This phenomenon and the vertical drift are caused by errors that happen during the acquisition and processing of scans. Although the average acquisition time of the OCT scans is 0.1 seconds, there are cases where the acquisition can take longer, causing a delayed update of the velocity, or shorter, causing multiple updates in a short time. Furthermore, wrongly segmented scans can lead to instances where the robot can receive a couple of velocity updates in the opposite direction followed by continued correct velocity updates, causing noisy or horizontal position lines. Accumulating the incorrect velocity changes leads to the drift seen in the results.
@@
Figure 7: OCT B-scan image captured at the 20-second mark of the insertion shown in Figure 6. A yellow circle marks the needle tip, the needle is highlighted in red, and the ILM layer of the retina is highlighted in green.
Another observation is the altered retina motion associated with the needle insertion. Looking at Fig. 6, the needle leaves the tissue by around 100 µm when the needle and linear stage motions are out of sync. However, when looking at the OCT scans, we realize that the needle tip remains in the retina and the horizontal shift is much less significant. When the needle penetrates the retina, it likely induces surface adhesion and resistance within the tissue, restricting free movement. This effect is particularly evident during downward movements, where the needle’s presence within the tissue exerts a tethering force that impedes motion. In Fig. 7, we show an OCT B-scan at the 20-second mark where the needle tip is still in contact with the retina. According to the graph, the needle should be outside of the retinal tissue at this point. Therefore, when tool-tissue interactions are present, the linear stage controller’s position does not fully represent the position or motion.
@@ VI CONCLUSION
In this work, we presented our initial method and experiments for extending our previously developed autonomous robotic subretinal injection pipeline with motion compensation capabilities along the $Z$-axis. Based on real time acquisition and segmentation of small 3D volumes, called B5-scans, we calculate the depth of the ILM layer and compare it with the previously computed values. We then move the robot end-effector to keep the needle tip in the same position relative to the retina.
Our method shows the potential of our robotic system to compensate for simulated motion like motion caused by respiration or circulation. However, some aspects should be further examined, such as the interactions of moving tissue and surgical tools and the implementation of predictive methods for estimating retinal motion for more accurate compensation. These are potential future work areas.